72.) Solve 3/x – 1/6 = 1/x

93) Solve x + 5 = 6/x

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(57222)

72.) Solve 3/x – 1/6 = 1/x
Multiply thru by 6x to get:
6*3 - x = 6
x = 18-6
x = 12
------------------
93) Solve x + 5 = 6/x
Multiply thru by x to get:
x^2 + 5x - 6 = 0
(x+6)(x-1) = 0
x = -6 or x = 1
====================
Cheers,
Stan H.

Answer by solver91311(16868)




LCD is 6x:



Multiply both sides by 6x:







Check:





Answer checks.

=======================================



Multiply both sides by x, then add -6 to both sides:



Factor:



So:

or

Check:

Checks.


Checks.